### 13 Mar 2026 v1

# Critical Sections Are Not Per-Thread: A Trace Semantics for Lock-Based Concurrency

Abstract.
Locks are a standard mechanism for synchronizing concurrent threads. The standard lock set construction assumes that critical sections are confined to a single thread, and therefore only accounts for locks acquired within that thread. The commonly used notion of a critical section implicitly assumes that protected events belong to the same thread. We show that this assumption is not valid for general C/Pthread executions.
Using a trace model that captures the essence of C/Pthread programs, we give a trace-based characterization of critical sections that does not impose a per-thread restriction. As a result, critical sections may span multiple threads. Such multi-thread critical sections arise naturally in real programs and close a semantic gap in the standard lock set construction.
1. Introduction
1#include <pthread.h>
2pthread_mutex_t m1 = PTHREAD_MUTEX_INITIALIZER;
3pthread_mutex_t m2 = PTHREAD_MUTEX_INITIALIZER;
4void* thread_m1_m2(void*) {
5 pthread_mutex_lock(&m1);
6 pthread_mutex_lock(&m2);
7 pthread_mutex_unlock(&m2);
8 pthread_mutex_unlock(&m1);
9return NULL;
10}
11void* thread_m1(void*) {
12 pthread_mutex_lock(&m1);
13 pthread_mutex_unlock(&m1);
14return NULL;
15}
16int main() {
17 pthread_t tid2, tid3;
18 pthread_create(&tid2, 0, thread_m1_m2, 0);
19 pthread_mutex_lock(&m2);
20 pthread_create(&tid3, 0, thread_m1, 0);
21 pthread_join(tid2, 0);
22 pthread_mutex_unlock(&m2);
23}
Figure 1. C/Pthread program.
We consider concurrent programs that make use of threads and locks. Many program analyses for deadlock and data race detection rely on the notion of a critical section, which informally describes the region of execution protected by a lock. In the literature, critical sections are typically defined as per-thread regions delimited by a lock acquisition and the corresponding release in the same thread. This definition implicitly assumes that the events protected by a lock belong to the same thread as the lock acquisition. However, this assumption does not hold for general executions of C/Pthread programs.
For example, consider the program in Figure 1. We assume that threads are numbered where thread 1 is the thread in which the main function executes. At line 18 another thread 2 is created. Thread 2 acquires lock m1 and then acquires lock m2 before releasing the locks. We can argue that the lock acquire operation at line 6 is part of the critical section with entry point at line 5 and exit point at line 8.
Entry/exit points of critical sections are pairs of lock/unlock operations that operate on the same lock variable and belong to the same thread. Anything that must occur between the entry and exit points in all execution traces belongs to the critical section. The common assumption is that ‘must inbetween’ coincides with thread order, and therefore that critical sections are per-thread. This assumption is implicit and fundamental to the standard lock set construction [DS91].
The above implicit assumptions do not apply to all C/Pthread programs as shown by our example in Figure 1. The thread creation operation at line 18 and the join operation at line 21 guarantee that the lock acquire operation at line 12 is always in between the critical section represented by the entry point at line 19 and the exit point at line 22. However, entry/exit points belong to a distinct thread. We conclude that the common notion of a critical section found in the literature is semantically incomplete with respect to execution traces of C/Pthread programs. Therefore, the standard lock set construction that computes the set of locks that are held (acquired) but not yet released at a certain point in the execution is incomplete as well.
Contributions
The contributions of this paper are primarily conceptual:
  • We identify an implicit per-thread assumption in the standard definition of critical sections and lock sets, and show that it is semantically incomplete with respect to a trace model that captures the essence of C/Pthread programs.
  • We give the first trace-based semantics of critical sections and lock sets that is complete for a minimal trace model capturing C/Pthread execution traces. The semantics captures multi-thread critical sections in execution traces.
Outline
Section 2 introduces a minimal trace model for C/Pthread programs. Section 3 reviews the commonly used per-thread definition of a critical section and lock set and present our more general trace-based characterization of a critical section and lock sets Section 4 discusses earlier works that rely on critical sections and lock sets for program analysis purposes. Section 5 concludes.
2. Events and Traces
We describe the semantics of programs in terms of execution traces. A trace is a list of events reflecting the interleaved execution of a concurrent program where each event is connected to some concurrency primitive. We consider a trace model with events to represent lock/unlock operations on locks and fork/join operations on threads. This minimal model captures the essence of C/Pthread programs.
{defi}
[Events and Traces]
$\begin{array}[]{@{}l@{}c@{}lll@{}c@{}llr@{}}e&{}::={}&(\alpha,t,op)&\text{(events)}\\ \alpha,\beta,\delta&{}::={}&1\mid 2\mid\ldots&\text{(unique event ids)}\\ \tau,\upsilon&{}::={}&\tau_{1}\mid\tau_{2}\mid\ldots&\text{(thread ids)}\\ op&{}::={}&\mathtt{fork}(\tau)\mid\mathtt{unlock}(m)\mid\mathtt{join}(\tau)\mid\mathtt{lock}(m)&\text{(operations)}\\ T&{}::={}&[]\mid e:T&\text{(traces)}\end{array}$
We write $o_{1},\dots,o_{n}]$ for a list of objects as a shorthand of $o_{1}:\dots:o_{n}:[]$ and use the operator $\mathbin{\cdot}{}$” for list concatenation.
An event $e$ is represented by a triple $(\alpha,\tau,op)$ where $\alpha$ is a unique event identifier, $op$ is an operation, and $\tau$ is the thread id in which the operation took place. Via the unique event identifier we can unambiguously distinguish among events that occur in the same thread and result from the same operation. For brevity, we often omit the event identifier when denoting events and write $e=(\tau,op)$ instead of $e=(\alpha,\tau,op)$. We write $\mathtt{fork}(\tau)$ for the creation of a new thread with thread id $\tau$ and $\mathtt{join}(\tau)$ for a join with a thread with thread id $\tau$. We write $\mathtt{lock}(m)$ and $\mathtt{unlock}(m)$ for locking and unlocking of some lock $m$.
[…]
Figure 2. Trace $T_{1}$ resulting from execution of the program in Figure 1.
{exa}
Figure 2 shows some trace $T_{1}$ resulting from execution of the program in Figure 1. We assume that the main thread always has thread id $\tau_{1}$. For ease of reading, we use a tabular notation for traces where each row contains exactly one event, in the column of the executing thread. The textual order (from top to bottom) reflects the observed temporal order of events. We write $e_{i}$ to refer to the $i$th in $T_{1}$. For example, $e_{5}=(5,\tau_{2},\mathtt{unlock}(m_{1}))$. In more compact form, trace $T_{1}$ can be written as $e_{1},...,e_{11}]$.
Traces are well formed and respect the program’s semantics. For example, a lock cannot be acquired twice etc. In the below, we write $l,l^{\prime}$ to refer to lock events and $u,u^{\prime}$ to refer to unlock events. We write $e\in T$ to indicate that $T=[e_{1},...,e_{n}]$ where $e=e_{k}$ for some $k$, defining also that $\mathit{pos}_{T}(e)=k$. For events $e,f\in T$ we define trace order: $e<_{tr}^{T}f$ if $\mathit{pos}_{T}(e)<\mathit{pos}_{T}(f)$.
{defi}
[Well Formedness] Trace $T$ is well formed if all the following conditions are satisfied:
WF-Acq:
For all $l=(\tau,\mathtt{lock}(m)),l^{\prime}=(\tau^{\prime},\mathtt{lock}(m))\in T$ where $l<_{tr}^{T}l^{\prime}$, there exists $u=(\tau,\mathtt{unlock}(m))\in T$ such that $l<_{tr}^{T}u<_{tr}^{T}l^{\prime}$.
WF-Rel:
For all $u=(\tau,\mathtt{unlock}(m))\in T$ there exists $l=(\tau,\mathtt{lock}(m))\in T$ such that $l<_{tr}^{T}u$ and there is no $u^{\prime}=(\tau^{\prime},\mathtt{unlock}(m))\in T$ with $l<_{tr}^{T}u^{\prime}<_{tr}^{T}u$.
WF-Fork1:
For all $\tau\neq\tau_{1}$ there exists at most one $e=(\upsilon,\mathtt{fork}(t))\in T$ and for all $e=(\upsilon,\mathtt{fork}(\tau))\in T$ we have that $\tau\neq\tau_{1}$.
WF-Fork2:
For all $e=(\tau,op)\in T$ where $\tau\neq\tau_{1}$ there exists $f=(\upsilon,\mathtt{fork}(\tau))\in T$ where $f<_{tr}^{T}e$.
WF-Join1:
For all $e=(\upsilon,\mathtt{join}(\tau))\in T$ we have $\upsilon\neq\tau$ and there exists $f=(\tau,op)\in T$.
WF-Join2:
For all $e=(\upsilon,\mathtt{join}(\tau))\in T$ and all $f=(\tau,op)\in T$ we have $f<_{tr}^{e}$.
WF-Acq states that a previously acquired lock can only be acquired after it has been released. Similarly, WF-Rel states that a lock can only be released after it has been acquired but not yet released. Both conditions require that the operation that releases a lock must occur in the same thread as the operation that acquired the lock.
WF-Fork1 states that a thread can be created at most once. WF-Fork2 states that each thread except the main thread is preceded by a fork event. WF-Join1 states that a join operation $\mathtt{join}(t)$ must occur in a thread distinct from $t$ and there must be operations that are executed in thread $t$. WF-Join2 states that all events from a joined thread must appear before the join event.
{exa}
Trace $T_{1}$ is well-formed. Consider the following traces where $e_{i}$ refer to the events in Figure 2:
$\begin{array}[]{lcl}T_{2}&=&[e_{1},e_{6},e_{7},e_{8},e_{9},e_{10},e_{11},e_{2},e_{3},e_{4},e_{5}]\\ T_{3}&=&[e_{1},e_{6},e_{7},e_{8},e_{9},e_{10},e_{11},e_{2}]\\ T_{4}&=&[e_{1},e_{7},e_{8},e_{10},e_{6},e_{11}]\\ T_{5}&=&[e_{1},e_{2},e_{6}]\\ T_{6}&=&[e_{2},e_{3},e_{4},e_{5}]\par\end{array}$
Traces $T_{2}$, $T_{3}$ and $T_{4}$ are well-formed as well. Traces $T_{5}$ and $T_{6}$ are ill-formed. Trace $T_{5}$ violates WF-Acq and trace $T_{6}$ violates WF-Fork2.
A trace represents one possible interleaving of concurrent events. In theory, there can be as many interleavings as there are (well-formed) reorderings of the original trace. However, not all trace reorderings are feasible in the sense that they could be reproduced by executing the program with a different schedule.
{exa}
Traces $T_{2}$, $T_{3}$ and $T_{4}$ are well-formed reorderings of a subset of the events in $T_{1}$. Traces $T_{2}$ and $T_{3}$ represent executions where a different schedule is taken compared to $T_{1}$. In case of trace $T_{2}$, the operations in thread $\tau_{2}$ are executed last. Trace $T_{3}$ is a variant of $T_{2}$ where we assume that not all operations in thread $\tau_{2}$ are executed because the main thread $\tau_{1}$ terminates immediately after $e_{11}$ and thus all other threads are terminated as well. Trace $T_{4}$ however results from completely different program run because the order among events in a thread has changed. In fact, we can argue there is no program run of the example in Figure 1 that leads to trace $T_{4}$.
We characterize all trace reorderings of trace $T$ that represent alternative executions under a different schedule in terms of correctly reordered prefixes. Reorderings must be correct in the sense that the reordered trace is well-formed and maintains order among events per thread as in the original trace $T$. Reorderings do not have to cover all events, so we consider reordered prefixes of traces. We make use of the following helper functions and definitions.
The projection of $T$ onto thread $\tau$, denoted $\mathit{proj}(T,\tau)$, restricts $T$ to events $e$ with where $e=(\tau,op)$. We say trace $T$ prefixes trace $T^{\prime}$ if $T^{\prime}=T\mathbin{\cdot}T^{\prime\prime}$ for some trace $T^{\prime\prime}$. We define $\mathit{thds}(T)=\{\tau\mid\exists e=(\tau,op)\in T\}$. For example, $\mathit{proj}(T_{3},\tau_{2})=[e_{2}]$ and $\mathit{evts}(T_{3})=\{e_{1},e_{6},e_{7},e_{8},e_{9},e_{10},e_{11},e_{2}\}$ and $\mathit{thds}(T_{1})=\{\tau_{1},\tau_{2},\tau_{3}\}$.
{defi}
[Correctly Reordered Prefix] Let $T$ be a well-formed trace. Then, trace  $T^{\prime}$ is a correctly reordered prefix of $T$ if the following conditions are satisfied:
CRP-WF:
$T^{\prime}$ is well formed.
CRP-PO:
For every $\tau\in\mathit{thds}(T^{\prime})$, $\mathit{proj}(T^{\prime},\tau)$ prefixes $\mathit{proj}(T,\tau)$.
We write $\mathit{crp}(T)$ to denote the set of correctly reordered prefixes of $T$.
CRP-PO states that the order of events in a thread must be maintained. CRP-WF states well-formedness and thus guarantees the conditions in Definition 2. For example, $T_{2},T_{3}\in\mathit{crp}(T_{1})$ and $T_{4},T_{5},T_{6}\not\in\mathit{crp}(T_{1})$.
3. Critical Section and Lock Set
We consider a specific program run represented by some well-formed trace $T$. In trace $T$ we wish to identify critical sections that protect events. The entry and exit points of a critical section are easily identified by matching lock and unlock events.
{defi}
[Matching Lock and Unlock Events] Let $T$ be a well-formed trace with events $l=(t,\mathtt{lock}(m)),u=(t,\mathtt{unlock}(m))\in T$. We say that $l$ and $u$ are matching lock and unlock events, written $\mathit{match}(l,u)_{T}$, if $l<_{tr}^{T}u$ and there exists no $u^{\prime}=(t,\mathtt{unlock}(m))\in T$ such that $l<_{tr}^{T}u^{\prime}<_{tr}^{T}u$. Due to well-formedness, each lock event has its distinct matching unlock event and vice versa. To avoid special treatment of ’open’ critical sections where the unlock event is missing we impose the following additional condition on traces:
WF-Acq2:
For all $l=(t,\mathtt{lock}(m))\in T$ there exists $u=(t,\mathtt{unlock}(m))\in T$ where $l<_{tr}^{T}u$.
The extra condition WF-Acq2 imposes no restriction as we can always add missing unlock events at the end of a trace.
Informally, an event is part of a critical section if it is enclosed by a lock event and its matching unlock event. In such a situation we say that the event holds the lock. The set of locks held is commonly referred to as the lock set. We first review the existing per-thread definition of a critical section and the resulting lock set construction and then introduce our more general trace-based characterization.
3.1. Per-Thread Definition
The following definition of a critical section and lock set is commonly employed in the literature.
{defi}
[Per-Thread Critical Section] Let $T$ be a well-formed trace with events $l=(\tau,\mathtt{lock}(m)),u=(\tau,\mathtt{unlock}(m))\in T$ where $\mathit{match}(l,u)_{T}$. We say that $e\in T$ is in the per-thread critical section guarded by $l$ and $u$ in thread $\tau$, written $e\in CS_{T}^{\tau}(l,u)$, if $l<_{tr}^{T}e<_{tr}^{T}u$ and $e=(\alpha,\tau,op)$ for some operation in the same thread as $l$ and $u$.
{defi}
[Per-Thread Locks Held] Let $T$ be a well-formed trace with an event $e=(\tau,op)\in T$. The per-thread lock set for an event $e\in T$ in thread $\tau$ is defined by $LH_{T}^{\tau}(e)=\{m\mid\exists l=(\tau,\mathtt{lock}(m)),u=(\tau,\mathtt{unlock}(m))\in T.e\in CS_{T}^{\tau}(l,u)\}$.
The superscript $\tau$ in $CS_{T}^{\tau}(l,u)$ and $LH_{T}^{\tau}(e)$ highlights that the above definitions are restricted to events that are in the same thread $\tau$.
{exa}
For trace $T_{1}$ in Figure 2 and events $e_{3}=(\tau_{2},\mathtt{lock}(m_{2})),e_{8}=(\tau_{3},\mathtt{lock}(m_{2}))\in T_{1}$ we find that $LH_{T_{1}}^{\tau_{2}}(e_{3})=\{m_{1}\}$ and $LH_{T_{1}}^{\tau_{3}}(e_{8})=\{\}$.
We can show that the lock set is stable in case program executions follows a different schedule of events.
Lemma 1.
Let $T$ be a well-formed trace with some event $e=(\tau,op)\in T$ where $m\in LH_{T}^{\tau}(e)$. Then, $m\in LH_{T^{\prime}}^{\tau}(e)$ for any reordering $T^{\prime}\in\mathit{crp}(T)$ where $e\in T^{\prime}$ and $T^{\prime}$ additionally satisfies WF-Acq2.
Proof 3.1.
By assumption, there exists $l=(\tau,\mathtt{lock}(m)),u=(\tau,\mathtt{unlock}(m))\in T$ such that $\mathit{match}(l,u)_{T}$ and $l<_{tr}^{T}e<_{tr}^{T}u$. Each $T^{\prime}\in\mathit{crp}(T)$ maintains the program order condition CRP-PO and with the additional condition WF-Acq2 we find that $l<_{tr}^{T^{\prime}}e<_{tr}^{T^{\prime}}u$. Hence, $e\in CS_{T^{\prime}}^{\tau}(l,u)$ and thus $m\in LH_{T^{\prime}}^{\tau}(e)$.
However, the per-thread lock set definition is semantically incomplete. For the running example in Figure 2 we find that for all reorderings $T^{\prime}\in\mathit{crp}(T)$ that additionally satisfy WF-Acq2 we have that event $e_{8}$ is surrounded by events $e_{6}$ and $e_{11}$. That is, $e_{6}<_{tr}^{T^{\prime}}e_{8}<_{tr}^{T^{\prime}}e_{11}$. This means that based on the observed execution represented by trace $T_{1}$, event $e_{8}$ is always protected by lock $m_{1}$ not matter which alternative schedule of events is taken. Due to the per-thread restriction of the lock set $LH_{T_{1}}^{\tau_{3}}(e_{8})$ lock $m_{1}$ is not an element in this set.
3.2. Trace-Based Characterization
To guarantee that the lock set is semantically complete, we give a trace-based characterization that lifts the per-thread restriction. For this purpose, we introduce the following ordering relation. We say event $f$ must be preceded by event $e$ under all trace reorderings, written $e<_{crp}^{T}f$, if $\forall T^{\prime}\in\mathit{crp}(T)$ where $f\in T^{\prime}$ we have that $e\in T^{\prime}$ and $e<_{tr}^{T^{\prime}}f$.
{defi}
[Critical Section] Let $T$ be a well-formed trace with events $l=(\tau,\mathtt{lock}(m)),u=(\tau,\mathtt{unlock}(m))\in T$ where $\mathit{match}(l,u)_{T}$. We say that $e\in T$ is in the critical section guarded by $l$ and $u$, written $e\in CS_{T}(l,u)$, if $l<_{crp}^{T}e<_{crp}^{T}u$.
The definition of a critical section is no longer restricted to a single thread. A critical section may cover multiple threads as can be seen by the example in Figure 2.
{defi}
[Locks Held] Let $T$ be a well-formed trace with an event $e=(\tau,op)\in T$. The lock set for an event $e\in T$ in thread $\tau$ is defined by $LH_{T}(e)=\{m\mid\exists l=(\tau,\mathtt{lock}(m)),u=(\tau,\mathtt{unlock}(m))\in T.e\in CS_{T}(l,u)\}$.
Lemma 2.
Let $T$ be a well-formed trace with an event $e=(\tau,op)\in T$. Then, $LH_{T}^{\tau}(e)\subseteq LH_{T}(e)$.
The subset relation is strict as shown by the following example.
{exa}
For trace $T_{1}$ in Figure 2 and events $e_{3}=(\tau_{2},\mathtt{lock}(m_{2})),e_{8}=(\tau_{3},\mathtt{lock}(m_{2}))\in T_{1}$ we find that $LH_{T_{1}}(e_{3})=\{m_{1}\}$ and $LH_{T_{1}}(e_{8})=\{m_{1}\}$.
Thanks to the lifted per-thread restriction, lock set is now semantically complete. We say that lock $m$ protects event $e\in T$ if there exist $l=(\tau,\mathtt{lock}(m)),u=(\tau,\mathtt{unlock}(m))\in T$ where $\mathit{match}(l,u)_{T}$ and for all $T^{\prime}\in\mathit{crp}(T)$ where $l,u\in T^{\prime}$ we have that $l<_{crp}^{T^{\prime}}e<_{crp}^{T^{\prime}}u$. Any lock $m$ that protects some event is in the event’s lock set. The following result follows by definition.
Lemma 3 (Semantic Completeness of Trace-Based Lock Set).
Let $T$ be a well-formed trace and $e\in T$. If lock $m$ protects $e$ then $m\in LH_{T}(e)$.
We can also restate lock set stability.
Lemma 4 (Lock Set Stability).
Let $T$ be a well-formed trace with some event $e=(\tau,op)\in T$ where $m\in LH_{T}(e)$. Then, $m\in LH_{T^{\prime}}(e)$ for any reordering $T^{\prime}\in\mathit{crp}(T)$ where $e\in T^{\prime}$ and $T^{\prime}$ additionally satisfies WF-Acq2.
4. Related Work
The notion of a critical section goes back to Dijkstra [Dij65]. It is hard to find a mathematical definition. We review earlier works in the context of program analysis that rely on the notion of critical section and lock set.
Deadlock detection
The use of lock set has emerged in the context of deadlock detection. For example, see the earlier dynamic deadlock detection works by Havelund [Hav00] and Harrow [Har00]. There are numerous follow-up works, e.g., see [BH05, JPSN09, CMP20, SR14, ZSL+17, CC14, KP18, CYW+21, TMPV23].
Our results apply independently of how traces are obtained (e.g., via instrumentation, symbolic execution, or CFG reachability). Static analysis works typically consider the reachability graph. There are numerous works that employ the lock set method for static deadlock detection, e.g., see [KPSW16, Sha08, NPSG09, NPSG09, CYSZ22, KPSW16].
Data race detection
The notion of critical section and lock set is also employed by works that cover data race detection. For example, consider lock set based data race method [SBN+97, SI09, XXZ13, YB16] and works that order conflicting critical sections [SES+12, KMV17, RGB18, GRXB19, SS20].
In all these works, critical sections and lock sets are implicitly defined per thread.
5. Conclusion
Pretty much all prior works assume per-thread lock sets and critical sections restricted to a single thread. This assumption does not hold for general C/Pthread executions. Critical sections that cover multiple threads exist, can be observed in simple programs and invalidate the standard foundation. Similar observations apply to Java as well. Correcting this semantic gap is important because lock sets form the foundation of numerous analyses for deadlock and data race detection. We consider a minimal trace model for C/Pthread programs and give corrected definitions based on a trace-based characterization of critical sections and lock sets to deal with programs where locks protect events across thread boundaries. Extending the semantics with memory operations and additional synchronization primitives such as condition variables introduces orthogonal challenges and is left for future work.